# Quantum speed limits based on Jensen-Shannon and Jeffreys divergences for general physical processes

### v1

@@ V.2.3 Generalized amplitude damping channel
In Fig. 5, we show the density plots of the entropic QSL time $\tau_{JS}^{\text{QSL}}$}, as a function of the dimensionless parameters $\Gamma\tau$ and $r$. We note that, for a fixed value $r\in[0,1)$ of the mixedness parameter, $\tau_{JS}^{\text{QSL}}$} varies monotonically as a function of the dimensionless parameter $\Gamma\tau\geq 0$. Overall, the entropic QSL approaches small values at earlier times of the dynamics, exhibiting similar qualitative behavior regardless the initial state and temperature of the thermal bath that we consider here. In particular, for $\alpha=\{0,1\}$ and $\theta=\pi/2$, Figs. 5(g) and 5(i) show that the QSL time exhibits the same quantitative behavior. This comes from the fact that $\tau_{JS}^{\text{QSL}}$} reduces to the same result under this choice of parameters, regardless of $r\in[0,1)$ and $\Gamma\tau\geq 0$ [see Eq. (27)]. This means that, for a given probe mixed state lying in the equatorial plane of the Bloch sphere, with $0<r<1$ and $\theta=\pi/2$, the QSL time at which the system reaches the stationary states $|1\rangle\langle{1}|$ ($\alpha=0$) or $|0\rangle\langle{0}|$ ($\alpha=1$) will be the same.
@@
In Fig. 6, we show the density plots of the normalized relative error ${\widetilde{\delta}_{JS}}({\tau})$, as a function of the dimensionless parameters $\Gamma\tau$ and $r$. We note that, for a fixed value $r\in[0,1)$ of the mixedness parameter, ${\widetilde{\delta}_{JS}}({\tau})$ monotonically decreases as a function of the dimensionless parameter $\Gamma\tau\geq 0$. In particular, for $\Gamma\tau\lesssim 0.25$, one finds that $0.8\lesssim{\widetilde{\delta}_{JS}}({\tau})\lesssim 1$, with $r\in[0,1)$. This means that the entropic QSL bound is looser at earlier times of the dynamics, regardless of the purity of the initial single-qubit state. Furthermore, we have that ${\widetilde{\delta}_{JS}}({\tau})$ approaches small values at later times of the dynamics. For example, Figs. 6(c), 6(f), 6(g), 6(h), and 6(i) show that ${\widetilde{\delta}_{JS}}({\tau})\lesssim 0.2$ for $0\leq r\lesssim 0.2$ and $\Gamma\tau\gtrsim 2$, with $\theta=\{0,\pi/4,\pi/2\}$ and $\alpha=\{0,0.1,1\}$.
@@ VI Conclusions
To illustrate our findings, we have specialized these results to the case of two-level systems undergoing (i) unitary dynamics [see Sec. V.1], and (ii) nonunitary dynamics [see Sec. V.2]. On the one hand, for closed quantum systems, we consider the unitary evolution of mixed single qubit generated by a time-independent Hamiltonian. We obtain analytical QSLs based on QJPD [see Eq. (30)] and QJSD [see Eq. (31)], also discussing their dynamical properties. On the one hand, for open quantum systems, we set the nonunitary dynamics generated by three quantum operations, namely, depolarizing channel, phase damping channel, and generalized amplitude damping channel. We provide analytical results for QJPD and QJSD for each case, and also present numerical simulations to support our findings on the QSLs [see Figs. 123, and 5]. We also address the tightness of these speed limits [see Figs. 124, and 6]. The bounds are looser at earlier times of the dynamics, and they are not tight for the overall evolution.
It is worth mentioning that entropic QSLs based on $\alpha$-$z$-Rényi relative entropies were recently presented in Ref. de Sousa and Pires (2025). Hence, it is natural to ask in what extent such results can or can not overcome the bounds discussed here, particularly since it would be possible to recover Umegaki’s relative entropy from $\alpha$-$z$-Rényi relative entropies by choosing $z=1$, and taking the limit $\alpha\rightarrow 1$. We point out that the results discussed here are not fully compatible with those addressed in Ref. de Sousa and Pires (2025). First, the results in Ref. de Sousa and Pires (2025) does apply only for $\alpha\in(0,1)$ and $1\geq z\geq\text{max}\{\alpha,1-\alpha\}$, with $\alpha\neq 1$, which rules out any chance of obtaining QSLs based on standard relative entropy from $\alpha$-$z$-Rényi relative entropies. Second, the number of inequalities used can actually lead to different families of bounds, which also undermines the tightness of these quantities. Indeed, tighter bounds on QJPD and QJSD are expected by applying a small number of inequalities, which may imply tighter and attainable speed limits. Leveraging the sensitivity of our QSLs respective to a few eigenvalues of the system overcomes the need to evaluate the full spectrum of effective Liouvillians in open quantum systems, which is the case of Refs. Mohan et al. (2022); Pandey et al. (2024, 2023) that showed QSLs based on relative entanglement entropies.

### v2

@@ V.2.3 Generalized amplitude damping channel
In Fig. 6, we show the density plots of the entropic QSL time $\tau_{JS}^{\text{QSL}}$}, as a function of the dimensionless parameters $\Gamma\tau$ and $r$. We note that, for a fixed value $r\in[0,1)$ of the mixedness parameter, $\tau_{JS}^{\text{QSL}}$} varies monotonically as a function of the dimensionless parameter $\Gamma\tau\geq 0$. Overall, the entropic QSL approaches small values at earlier times of the dynamics, exhibiting similar qualitative behavior regardless the initial state and temperature of the thermal bath that we consider here. In particular, for $\alpha=\{0,1\}$ and $\theta=\pi/2$, Figs. 6(g) and 6(i) show that the QSL time exhibits the same quantitative behavior. This comes from the fact that $\tau_{JS}^{\text{QSL}}$} reduces to the same result under this choice of parameters, regardless of $r\in[0,1)$ and $\Gamma\tau\geq 0$ [see Eq. (27)]. This means that, for a given probe mixed state lying in the equatorial plane of the Bloch sphere, with $0<r<1$ and $\theta=\pi/2$, the QSL time at which the system reaches the stationary states $|1\rangle\langle{1}|$ ($\alpha=0$) or $|0\rangle\langle{0}|$ ($\alpha=1$) will be the same.
@@
In Fig. 7, we show the density plots of the normalized relative error ${\widetilde{\delta}_{JS}}({\tau})$, as a function of the dimensionless parameters $\Gamma\tau$ and $r$. We note that, for a fixed value $r\in[0,1)$ of the mixedness parameter, ${\widetilde{\delta}_{JS}}({\tau})$ monotonically decreases as a function of the dimensionless parameter $\Gamma\tau\geq 0$. In particular, for $\Gamma\tau\lesssim 0.25$, one finds that $0.8\lesssim{\widetilde{\delta}_{JS}}({\tau})\lesssim 1$, with $r\in[0,1)$. This means that the entropic QSL bound is looser at earlier times of the dynamics, regardless of the purity of the initial single-qubit state. Furthermore, we have that ${\widetilde{\delta}_{JS}}({\tau})$ approaches small values at later times of the dynamics. For example, Figs. 7(c), 7(f), 7(g), 7(h), and 7(i) show that ${\widetilde{\delta}_{JS}}({\tau})\lesssim 0.2$ for $0\leq r\lesssim 0.2$ and $\Gamma\tau\gtrsim 2$, with $\theta=\{0,\pi/4,\pi/2\}$ and $\alpha=\{0,0.1,1\}$.
@@ VI Conclusions
To illustrate our findings, we have specialized these results to the case of two-level systems undergoing (i) unitary dynamics [see Sec. V.1], and (ii) nonunitary dynamics [see Sec. V.2]. On the one hand, for closed quantum systems, we consider the unitary evolution of mixed single-qubit states generated by a time-independent Hamiltonian. We obtain analytical QSLs based on QJPD [see Eq. (30)] and QJSD [see Eq. (31)], also discussing their dynamical properties. On the other hand, for open quantum systems, we set the nonunitary dynamics generated by three quantum operations, namely, depolarizing channel, phase damping channel, and generalized amplitude damping channel. We provide analytical results for QJPD and QJSD in each case, and also present numerical simulations to support our findings about QSLs [see Figs. 234, and 6]. We also address the tightness of these speed limits [see Figs. 235, and 7]. The bounds are looser at earlier times of the dynamics, and they are not tight for the overall evolution.
It is worth mentioning that entropic QSLs based on $\alpha$-$z$-Rényi relative entropies were recently presented in Ref. de Sousa and Pires (2025). Hence, it is natural to ask in what extent such results can or can not overcome the bounds discussed here, particularly since it would be possible to recover Umegaki’s relative entropy from $\alpha$-$z$-Rényi relative entropies by choosing $z=1$, and taking the limit $\alpha\rightarrow 1$. We point out that the results discussed here are not compatible with those addressed in Ref. de Sousa and Pires (2025). First, the results in Ref. de Sousa and Pires (2025) does apply only for $\alpha\in(0,1)$ and $1\geq z\geq\text{max}\{\alpha,1-\alpha\}$, with $\alpha\neq 1$, which rules out any chance of obtaining QSLs based on standard relative entropy from $\alpha$-$z$-Rényi relative entropies. This means that the bounds presented here cannot be recovered from the results in Ref. de Sousa and Pires (2025), neither as particular cases nor as straightforward applications. Second, the number of inequalities used can actually lead to different families of bounds, which also undermines the tightness of these quantities. Indeed, tighter bounds on QJPD and QJSD are expected by applying a small number of inequalities, which may imply tighter and attainable speed limits. Leveraging the sensitivity of our QSLs respective to a few eigenvalues of the system overcomes the need to evaluate the full spectrum of effective Liouvillians in open quantum systems, which is the case of Refs. Mohan et al. (2022); Pandey et al. (2024, 2023) that showed QSLs based on relative entanglement entropies.